If the half-life for polonium-218 is 3.05 minutes. How many half-lives have passed for polonium-218 after 9.15 minutes
How much of an 89.6mg sample polonium-218 would remain after 9.15 minutes
1
Expert's answer
2013-02-27T06:11:02-0500
The half-life period for Po-218is t_hl = 3.05 minutes. Therefore, after t = 9.15 minutes, it has been (t / t_hl) = (9.15 min / 3.05 min) = 3 half-life periods for this substance. If the initial mass was 89.6 mg of Po-218, after 9.15 minutes the remainder
would be
m = m_0 * (1/2)^(t/t_hl) = 89.6 mg * (0.5)^(9.15 min / 3.05 min) = 89.6 mg *
(1/8) = 11.2 mg.

Answer: 3 half-life periods; 11.2 mg.